### v1

@@ Abstract
We present an efficient method for adapting a monolingual Large Language Model (LLM) to another language, addressing challenges of catastrophic forgetting and tokenizer limitations. We focus this study on adapting Llama 2 to Arabic. Our two-stage approach begins with expanding the vocabulary and training only the embeddings matrix, followed by full model continual pre-training on a bilingual corpus. By continually pre-training on a mix of Arabic and English corpora, the model retains its proficiency in English while acquiring capabilities in Arabic. Our approach results in significant improvements in Arabic and slight enhancements in English, demonstrating cost-effective cross-lingual transfer. We also perform extensive ablations on embedding initialization techniques, data mix ratios, and learning rates and release a detailed training recipe.
@@ 1 Introduction
There has been a rapid advancement in open source English-dominant foundation language models like Llama 2 Touvron et al. (2023), Mistral 7B Jiang et al. (2023), and Llama 3, primarily trained on extensive English corpora with minimal inclusion of non-English languages. To create models proficient in low-resource languages, two approaches can be taken: training a bilingual or multilingual model from scratch or adapting an existing strong English-dominant model to the target language. While bilingual and monolingual models trained from scratch, like Jais Sengupta et al. (2023b) and Bloom et. al. (2023), have shown promise in non-English capabilities, they are expensive to train and have inferior capabilities in English. Adapting strong English-dominant models to new languages also pose challenges, such as catastrophic forgetting of English capabilities, inefficiencies of English-dominant tokenizers, and the need for hyperparameter adjustments. Fujii et al. (2024); Luo et al. (2024); French (1999); Huang et al. (2024). In this paper we focus on addressing the challenges of the model adaptation approach.
It has been shown that capabilities of Large Language Models (LLMs) such as knowledge, reasoning and truthfulness are transferable across languages Yang et al. (2024); Sengupta et al. (2023b). This gives us the basis and motivation to explore efficient methods for cross-lingual transfer from English to Arabic through continual pre-training of a monolingual English-dominant LLM without degradation of English capabilities. Several recent works demonstrate cross-lingual transfer of foundation models de Vries & Nissim (2021); Marchisio et al. (2023); Csaki et al. (2023); Zhao et al. (2024); Huang et al. (2024); Da Dalt et al. (2024), yet they lack comprehensive analysis of hyperparameter tuning, tokenizer and data mix selections, and the impact of different model sizes.
@@
Vocabulary extension We establish that adapting an existing model to a new language requires expanding the vocabulary, along with employing the methods below to maintain the model’s original capabilities while acquiring new linguistic skills. We determine the optimal extension ratio of the original vocabulary through experimentation.
Embedding alignment We find that it ensuring alignment between the embeddings of the original and newly added vocabulary tokens is vital. We explore three techniques for initializing newly added token embeddings. We follow with embedding-only pre-training, which further aligns the embedding scale and orientation for original and new tokens.
Continual pre-training Following the embedding-only pre-training, we unfreeze the transformer backbone to continually pre-train the full model. We conduct experiments at the 7B model scale to assess various English-Arabic mix ratios and learning rates. We leverage the insights obtained from these experiments to perform cross-lingual adaptation to Arabic with Llama 2 13B and Llama 2 70B models.
Careful experimental study of winning strategies for vocabulary extension, embedding alignment, Arabic and English data mixes and hyperparameter tuning, results in a recipe for language adaptation with significant performance improvements in Arabic and, uniquely, enhancements in English on Llama 2 models.
@@ 2 Pre-training Datasets
We use the AraV5 Arabic dataset, which includes documents from various sources such as web pages, Wikipedia, Arabic news outlets, books and social media. It also includes high-quality English corpora, Books and Wikipedia, translated to Arabic. Curated by Sengupta et al. (2023b), it was used for pre-training the Jais series of Arabic foundation models. Prior domain adaptation studies have emphasized the importance of using ”replay” data, which aligns with the pre-training data domain, to preserve the foundational model’s knowledge and reasoning capabilities, as proposed in works by Gupta et al. (2023), Chen et al. (2023), and Azerbayev et al. (2024). We use the Pile corpus, comprising data from 22 diverse sources including ArXiv, Wikipedia, PubmedCentral, CommonCrawl, OpenWebText, and Github Gao et al. (2020). For Llama 2 7B adaptation, we utilize 20B tokens from AraV5, whereas for Llama 2-13B and 70B, we utilize the entire AraV5 corpus, totaling 140B tokens.
@@ 3.1 Vocabulary Extension
This imbalance introduces inefficiency in pre-training, fine-tuning and inference. Table 1 shows that the Llama 2 tokenizer needs as many as times the number of tokens to represent the same Arabic text as Jais’ Arabic-English bilingual tokenizer (MLV2) Sengupta et al. (2023b). Balanced multilingual tokenizer offers three main advantages Petrov et al. (2023): 1) lower training and inference cost; 2) reduced latency during inference; 3) longer context windows.
We experiment with two methods, vocabulary replacement and vocabulary extension, to create balanced tokenizers for English and Arabic. Vocabulary replacement implies maintaining the base vocabulary and replacing its least frequent tokens with the most frequent Arabic tokens. Vocabulary extension adds the most frequent Arabic tokens, increasing the vocabulary size. In both methods, we ensure that the newly introduced tokens are not present in the original vocabulary. For both methods, we determine the optinmal number of new tokens to create a balanced multilingual vocabulary. Using Arabic tokens from the MLV2 vocabulary, we create two candidate tokenizers and perform intrinsic and extrinsic evaluations following Ali et al. (2024).
@@
We perform extrinsic evaluation by continually training Llama 2 7B on a mixture of AraV5 and Pile, and monitoring the AraV5 validation loss. For a fair comparison of the tokenizers, we fix the raw text bytes at GB for Pile and GB for AraV5. Using Llama 2-extend100 as the candidate tokenizer and Llama 2 as the baseline, we tokenize the raw corpora and continually pre-train Llama 2 7B. Although the base Llama 2 tokenizer achieves a lower AraV5 validation loss compared to Llama 2-extend100 (see Table 1), it is trained on significantly more Arabic tokens due to its times higher fertility in Arabic. In an iso-token comparison, where the number of AraV5 tokens is fixed, Llama 2-extend100 outperforms base Llama 2 tokenizer by . Considering both the intrinsic and extrinsic evaluations, we select Llama 2-extend100. We correct all losses to align with the Llama 2 tokenizer (see A and B).
@@ 3.2 Embedding initialization
For Llama-extend100, we add new Arabic tokens to the Llama 2 vocabulary, expanding the embedding and unembedding layers as where d is the hidden size of the transformer decoder block. Our studies reveal that a choice of embeddings initialization for newly added tokens is critical. Simple methods such as Xavier Glorot & Bengio (2010) or Kaiming He et al. (2015) initialization, or using the mean of the embedding vectors of Llama 2 tokens, do not yield satisfactory results. Therefore, we explore alternative methods, described below, which demonstrated superior performances.
@@
This method is inspired by the approach proposed in Minixhofer et al. (2022). For each new token, we identify the top similar tokens in the base vocabulary, using an external embedding. We use OpenAI’s text-embedding-3-large embeddings Kusupati et al. (2024) for their superior quality and multilingual performance. Using cosine similarity, we find top similar tokens a new token and initialize the new token embeddings by taking the weighted average of base embeddings of these similar tokens. After experimenting with different values for the , we achieve the best results with .
@@ 3.4 Hyper-parameter tuning
Hyperparameter sweep is important to determine best hyper-parameters such as learning rate (lr), warm-up schedule and batch size (bs). We use a linear warm-up for 1% of the total steps followed by cosine decay Loshchilov & Hutter (2017) to 1/10th of the peak learning rate. We compare batch sizes of 4M tokens and 6M tokens but don’t see a significant difference in upstream losses. We pick 4M tokens as the final batch size. Following Gupta et al. (2023), we setup a learning rate sweep taking three different learning rates in different ranges. Let be the peak Llama 2 learning rate which is , we “re-warm” the learning rate to i) , ii) , and iii) . These experiments use billion AraV5 tokens mixed with Pile in , and ratios. Across all ratios we find that performs the best as shown in table 2.
@@ 3.5 Data mixture
Domain adaptation involves continual pre-training a foundation model on new data not seen during the pre-training. When this new domain data is out-of-distribution, it can cause significant forgetting of prior capabilities. Adding a small proportion of general domain data, or replay data, can mitigate the forgetting. We conduct exhaustive experiments to find a minimum proportion of Pile data that should be mixed with AraV5 to mitigate forgetting. Table 2 shows results from the experiments with different data mixes. We found that mixing part English with parts Arabic ( En:Ar) is sufficient to mitigate forgetting. We also don’t see any forgetting in downstream evaluation as discussed in section 4. Interestingly, increasing the amount of English data while keeping Arabic tokens constant improves Arabic performance, indicating cross-lingual capability transfer.
@@ 4 Results
Using the methodology described in section 3 we adapt Llama 2 7B, 13B and 70B models to Arabic. We use linearly warm up of the learning rate to for the first % of the tokens followed by cosine decay to th of the . For 7B and 13B models, we use En:Ar mix as we show in section 3.5 that higher proportion of English also improves Arabic performance. For Llama 2 70B we use En:Ar mix to reduce compute time. Llama 2 7B adaptation uses billion tokens each from AraV5 and Pile. Llama 2-13B and Llama 2 70B use all billion tokens from AraV5 and billion and billion tokens from Pile, respectively. Tables 3 and 4 show the -shot downstream performance of the resulting models against the base models and other Arabic models like Jais Sengupta et al. (2023a) and AceGPT Huang et al. (2024). For Arabic evaluations we translated the English downstream tasks datasets using a similar approach as in Sengupta et al. (2023b).
We evaluated the models on the World Knowledge tasks MMLU Hendrycks et al. (2021) and Exams Hardalov et al. (2020); Commonsense reasoning tasks Hellaswag Zellers et al. (2019), PIQA Bisk et al. (2020), SIQA Sap et al. (2019), BoolQ Clark et al. (2019), Arc Challenge Clark et al. (2018) and OpenBookQA Mihaylov et al. (2018); Misinformation and Bias tasks TruthfulQA Lin et al. (2021) and CrowS-PairsNangia et al. (2020).
For all models we see improvement across all tasks in Arabic. We observe % improvement in Arabic MMLU for Llama2 70B adapted compared to Llama2 70B (see 3), while the smaller models (7B and 13B) demonstrate % improvement in MMLU Arabic. This can be attributed to the lower token-per-parameter training regime resulting in less degradation from over-training Hoffmann et al. (2022); Dey et al. (2023). We also observe slight improvement in average scores in English for Llama2 70B adapted (see4).
Furthermore, compared to the state-of-the-art Arabic models, namely Jais and AceGPT, Arabic adapted Llama2 models significantly improve on Arabic downstream tasks.
@@ 5 Conclusion
We present an efficient recipe to significantly enhance capabilities of an English-dominant foundational LLM in another language. Our approach includes extending the vocabulary, applying a novel method for embedding initialization and alignment, and continually pre-training the foundation LLM on a bilingual data mix. We perform hyper-parameter optimization for batch size, learning rate schedule, and data mix ratio to ensure successful adaptation without experiencing “catastrophic forgetting”. We successfully use this approach to enhance Arabic capability of Llama 2 base models, resulting in a state-of-the-art 70B Arabic base language model. Furthermore, we apply this approach for other languages such as Turkic and Hindi and other foundation LLMs, with results for these adaptations to be presented in the future.

### v2

@@ Abstract
We present an efficient method for adapting a monolingual Large Language Model (LLM) to another language, addressing challenges of catastrophic forgetting and tokenizer limitations. We focus this study on adapting Llama 2 to Arabic. Our two-stage approach begins with expanding the vocabulary and training only the embeddings matrix, followed by full model continual pre-training on a bilingual corpus. By continually pre-training on a mix of Arabic and English corpora, the model retains its proficiency in English while acquiring capabilities in Arabic. Our approach results in significant improvements in Arabic and slight enhancements in English, demonstrating cost-effective cross-lingual transfer. We perform ablations on embedding initialization techniques, data mix ratios, and learning rates and release a detailed training recipe. To demonstrate generalizability of this approach we also adapted Llama 3 8B to Arabic and Llama 2 13B to Hindi.
@@ 1 Introduction
There has been a rapid advancement in open source English-dominant foundation language models like Llama 2 Touvron et al. (2023), Mistral 7B Jiang et al. (2023), and Llama 3, primarily trained on extensive English corpora with minimal inclusion of non-English languages. To create models proficient in low-resource languages, two approaches can be taken: training a bilingual or multilingual model from scratch or adapting an existing strong English-dominant model to the target language. While bilingual and monolingual models trained from scratch, like Jais Sengupta et al. (2023b) and Bloom et. al. (2023), have shown promise in non-English capabilities, they are expensive to train and have inferior capabilities in English. Adapting strong English-dominant models to new languages also pose challenges, such as catastrophic forgetting of English capabilities, inefficiencies of English-dominant tokenizers, and the need for hyperparameter adjustments. Fujii et al. (2024); Luo et al. (2024); French (1999); Huang et al. (2024). Here we address the challenges of the model adaptation approach.
Knowledge, reasoning and truthfulness capabilities of Large Language Models (LLMs) are transferable across languages Yang et al. (2024); Sengupta et al. (2023b). This gives us the basis and motivation to explore efficient methods for cross-lingual transfer from English to Arabic through continual pre-training of a monolingual English-dominant LLM without degradation of English capabilities. Several recent works demonstrate cross-lingual transfer of foundation models de Vries & Nissim (2021); Marchisio et al. (2023); Csaki et al. (2023); Zhao et al. (2024); Huang et al. (2024); Da Dalt et al. (2024), yet they lack comprehensive analysis of hyperparameter tuning, tokenizer, data mix selections, and the impact of different model sizes.
@@
Vocabulary extension We find that adapting a pre-trained model to a new language requires expanding the vocabulary, along with employing the methods below to maintain the model’s original capabilities while acquiring new linguistic skills. We determine the optimal extension ratio of the original vocabulary through experimentation.
Embedding alignment We find that ensuring alignment between the embeddings of the original and newly added vocabulary tokens is vital. We explore three techniques for initializing newly added token embeddings. We follow with embedding-only pre-training, which further aligns the embedding scale and orientation for original and new tokens.
Continual pre-training Following the embedding alignment, we continually pre-train the model. We conduct experiments at the 7B model scale to assess various English-Arabic mix ratios and learning rates. We leverage the insights obtained from these experiments to perform cross-lingual adaptation to Arabic with Llama 2 13B and Llama 2 70B models.
Careful empirical study of vocabulary extension, embedding alignment, data mixtures and hyperparameter tuning, results in a recipe for language adaptation with significant performance improvements in Arabic and, uniquely, enhancements in English on Llama 2 models. To demonstrate the generalizability of this recipe to other languages and models, we adapt Llama 3 8B to Arabic and Llama 2 13B to Hindi.
@@ 2 Pre-training Datasets
We use the AraV5 Arabic dataset, which includes documents from various sources such as web pages, Wikipedia, Arabic news outlets, books and social media. It also includes high-quality English corpora, Books and Wikipedia, translated to Arabic. Curated by Sengupta et al. (2023b), it was used for pre-training the Jais series of Arabic foundation models. Prior domain adaptation studies have emphasized the importance of using ”replay” data, which aligns with the pre-training data domain, to preserve the foundational model’s knowledge and reasoning capabilities, as proposed in works by Gupta et al. (2023); Chen et al. (2023); Azerbayev et al. (2024). We use the Pile corpus, comprising data from 22 diverse sources including ArXiv, Wikipedia, PubmedCentral, CommonCrawl, OpenWebText, and Github Gao et al. (2020). For Llama 2 7B adaptation, we utilize 20B tokens from AraV5, whereas for Llama 2-13B and 70B, we utilize the entire AraV5 corpus, totaling 140B tokens. For Hindi adaptation of Llama 2 13B we use a Hindi corpus of 65B tokens curated from diverse resources and mix it with Pile.
@@ 3.1 Vocabulary Extension
This imbalance introduces inefficiency in pre-training, fine-tuning and inference. Table 1 shows that the Llama 2 tokenizer needs as many as times the number of tokens to represent the same Arabic text as Jais’ Arabic-English bilingual tokenizer (MLV2) Sengupta et al. (2023b). Balanced multilingual tokenizer with low fertility in all languages offers three main advantages Petrov et al. (2023): 1) lower training and inference cost; 2) reduced latency during inference; 3) longer context windows. Models trained with low fertility tokenizers tend to perform well on downstream tasks as shown in Ahuja et al. (2023)
We experiment with two methods, vocabulary replacement and vocabulary extension, to create balanced tokenizers for English and Arabic. Vocabulary replacement implies maintaining the base vocabulary and replacing its least frequent tokens with the most frequent Arabic tokens. Vocabulary extension adds the most frequent Arabic tokens, increasing the vocabulary size. In both methods, we ensure that the newly introduced tokens are not present in the original vocabulary. For both methods, we determine the optimal number of new tokens to create a balanced multilingual vocabulary. Using Arabic tokens from the MLV2 vocabulary, we create two candidate tokenizers and perform intrinsic and extrinsic evaluations following Ali et al. (2024).
@@
We perform extrinsic evaluation by continually training Llama 2 7B on a mixture of AraV5 and Pile, and monitoring the AraV5 validation loss. For a fair comparison of the tokenizers, we fix the raw text bytes at GB for Pile and GB for AraV5. Using Llama 2-extend100 as the candidate tokenizer and Llama 2 as the baseline, we tokenize the raw corpora and continually pre-train Llama 2 7B. Although the base Llama 2 tokenizer achieves a lower AraV5 validation loss compared to Llama 2-extend100 (see Table 1), it is trained on significantly more Arabic tokens due to its times higher fertility in Arabic. In an iso-token comparison, where the number of AraV5 tokens is fixed, Llama 2-extend100 outperforms base Llama 2 tokenizer by . Considering both the intrinsic and extrinsic evaluations, we select Llama 2-extend100. We correct all losses to align with the Llama 2 tokenizer (see A and B). We performed a similar analysis for Llama 3 8B adaptation to Arabic. Llama 3 8B has a four times larger vocabulary than Llama 2 and possesses superior multilingual capabilities including in Arabic. We found that extending it by 20% brought the Arabic fertility within 10% of the MLV2 tokenizer’s fertility rate for Arabic. For Hindi adaptation of Llama 2 we extended the base vocabulary by 100% just as for Llama 2 Arabic adaptation.
@@ 3.2 Embedding initialization
For Llama-extend100, we add new Arabic tokens to the Llama 2 vocabulary, expanding the embedding and unembedding layers as where d is the hidden size of the transformer decoder block. Our studies showed that the choice of embeddings initialization for newly added tokens is critical such that tokens which represent similar concepts in two languages are closer in the embeddings latent space. Simple methods such as Xavier Glorot & Bengio (2010) or Kaiming He et al. (2015) initialization, or using the mean of the embedding vectors of Llama 2 tokens, do not yield satisfactory results. Therefore, we explore alternative methods, described below, which demonstrated superior performances.
@@
This method is inspired by the approach proposed in Minixhofer et al. (2022). For each new token, we identify the top- similar tokens using cosine similarity in the base vocabulary, using an external embedding. We use OpenAI’s text-embedding-3-large embeddings Kusupati et al. (2024) for their superior quality and multilingual performance. Using cosine similarity, we find top- similar tokens for the new token and initialize the new token embeddings by taking the weighted average of base embeddings of these similar tokens. After experimenting with different values for the , we achieve the best results with . This initialization method was used for embeddings and unembeddings layers of Llama 2 13B adapted to Hindi.
@@ 3.4 Hyper-parameter tuning
Hyperparameter sweep is important to determine best hyper-parameters such as learning rate (lr), warm-up schedule and batch size (bs). We use a linear warm-up for 1% of the total steps followed by cosine decay Loshchilov & Hutter (2017) to 1/10th of the peak learning rate. We compare batch sizes of 4M tokens and 6M tokens but don’t see a significant difference in upstream losses. We pick 4M tokens as the final batch size. We setup a learning rate sweep taking three different learning rates in different ranges. Let be the peak Llama 2 learning rate which is , we “re-warm” the learning rate to i) , ii) , and iii) . These experiments use billion AraV5 tokens mixed with Pile in , and ratios. Across all ratios we find that performs the best as shown in table 2. We used same hyperparameters for Hindi adaptation of Llama 2 13B. For Llama 3 8B we performed a similar sweep and found that for stable adaptation to new domain. a smaller learning rate of , a larger batch size of 6M tokens is required.
@@ 3.5 Data mixture
Domain adaptation involves continual pre-training a foundation model on new data not seen during the pre-training. When this new domain data is out-of-distribution, it can cause significant forgetting of prior capabilities which is referred to as stability gap Guo et al. (2024). Adding a small proportion of replay data, which is closer in distribution to the pre-training data, can mitigate the forgetting. We conduct exhaustive experiments to find a minimum proportion of Pile data that should be mixed with AraV5 to mitigate forgetting. Table 2 shows results from the experiments with different data mixes. We found that mixing part English with parts Arabic ( En:Ar) is sufficient to mitigate forgetting. We also don’t see any forgetting in downstream evaluation as discussed in section 7. Interestingly, increasing the amount of English data while keeping Arabic tokens constant improves Arabic performance, indicating cross-lingual capability transfer. We leave the exploration of cross-lingual transfer to future work.
For Arabic adaptation of Llama 3 8B we found that a higher proportion of replay data is necessary.. Therefore, English to Arabic dataset mixture was used where we saw cross-lingual capability transfer between the two languages similar to Llama 2 adaptation. For replay data we needed a mix of textbooks, math, coding and reasoning datasets mitigate forgetting.
@@ 4 Results
| llama 3-8b base | 33.0% | 45.2% | 47.3% | 43.2% |
| llama 3-8b adapted | 38.8% | 49.2% | 53.3% | 47.9% |
| llama 3-8b base | 51.1% | 63.7% | 53.2% | 59.5% |
| llama 3-8b adapted | 49.9% | 64.0% | 58.9% | 60.5% |
| Model | XNLI Hindi | English downstream average |
| llama 2-13b base | 37,44% | 57.1% |
| llama 2-13b adapted | 45.68% | 55.3% |
Using the methodology described in section 3 we adapt Llama 2 7B, 13B and 70B models to Arabic. We use linear warm up of the learning rate to for the first % of the tokens followed by cosine decay to th of the . For 7B and 13B models, we use En:Ar mix as we show in section 3.5 that higher proportion of English also improves Arabic performance. For Llama 2 70B we use En:Ar mix to reduce compute time. Llama 2 7B adaptation uses billion tokens each from AraV5 and Pile. Llama 2-13B and Llama 2 70B use all billion tokens from AraV5 and billion and billion tokens from Pile, respectively. Tables 3 and 4 show the -shot downstream performance of the resulting models against the base models and other Arabic models like Jais Sengupta et al. (2023a) and AceGPT Huang et al. (2024). For Arabic evaluations we translated the English downstream tasks datasets using a similar approach as in Sengupta et al. (2023b).
We evaluated the models on the World Knowledge tasks MMLU Hendrycks et al. (2021) RACE Lai et al. (2017) and Exams Hardalov et al. (2020); Commonsense reasoning tasks Hellaswag Zellers et al. (2019), PIQA Bisk et al. (2020), SIQA Sap et al. (2019), BoolQ Clark et al. (2019), Arc Challenge Clark et al. (2018), OpenBookQA Mihaylov et al. (2018) and Winogrande Sakaguchi et al. (2019); Misinformation and Bias tasks TruthfulQA Lin et al. (2021) and CrowS-PairsNangia et al. (2020).
For all models we see improvement in Arabic tasks - % improvement in Arabic MMLU for Llama2 70B adapted compared to Llama2 70B (see 3), while the smaller models (7B and 13B) demonstrate % improvement in MMLU Arabic. This can be attributed to the lower token-per-parameter training regime resulting in less degradation from over-training Hoffmann et al. (2022); Dey et al. (2023). We also observe slight improvement in average scores in English for Llama2 70B adapted (see4). Furthermore, compared to the state-of-the-art Arabic models, namely Jais and AceGPT, Arabic adapted Llama2 models significantly improve on Arabic downstream tasks.
For Arabic adaptation Llama 3 8B was continually pre-trained for 65B tokens with mix of English and Arabic using hyper-parameters as discussed in 3.4. In 5 and 6 we depict the performance of Llama 3 8B adapted on Arabic and English tasks, respectively, and a comparison against the Llama 3 8B base model. We see % improvement on average in Arabic downstream tasks compared to the base model performance while a 1% improvement in downstream English tasks.
Llama 2 13B was trained on 72B tokens with English to Hindi dataset mixture. Learning rate was re-warmed to the pre-training peak learning rate of and a batch size of 4M tokens. We evaluate on XNLI-Hindi Conneau et al. (2018) where the goal is to predict textual entailment. We see % improvement compared to Llama 2 13B while in English there is % degradation as shown in 7.
@@ 5 Conclusion
We present an efficient recipe to significantly enhance capabilities of an English-dominant foundation LLM in another language. Our approach includes extending the vocabulary, applying a novel method for embedding initialization and alignment, and continually pre-training the foundation LLM on a bilingual data mix. We perform hyper-parameter optimization for batch size, learning rate schedule, and data mix ratio to ensure successful adaptation without experiencing “catastrophic forgetting”. We successfully use this approach to enhance Arabic capability of Llama 2 base models, resulting in a state-of-the-art 70B Arabic base language model. Furthermore, we apply this approach to other languages such as Hindi and other foundation LLMs such as Llama 3 - while we briefly discuss these two studies but will leave the details for future work. We also intend to extend this approach to low resource languages and architecture such as depth-up scaling Kim et al. (2024) and adapter layers methods as discussed in D.